# Next-to-Next-to-Leading Order QCD Corrections to Polarized Semi-Inclusive Deep-Inelastic Scattering

### v2

@@ Abstract
×$\times\mathcal{G}_{1,ab}\left(\frac{x}{x_{1}},\frac{z}{z_{1}},\mu_{F}^{2},Q^{2}\right)\,,$ (4)
@@
Beyond LO in perturbative QCD, we encounter both ultraviolet (UV) and infrared (IR) singularities. The latter are due to the presence of soft and collinear partons. We regulate these singularities using dimensional regularization with $D=4+\varepsilon$ space-time dimensions. The projection of spin-dependent partonic amplitudes squared $|\Delta M_{ab}|^{2}$ in eq. (6) requires Dirac matrices $\gamma_{5}$ or the Levi-Civita tensor for polarized quarks or gluons, respectively, see, e.g. Zijlstra and van Neerven (1994). Since $\gamma_{5}$ and the Levi-Civita tensor are intrinsically four-dimensional objects, their treatment in $D$ dimensions requires some prescription. Although, several schemes to define them in $D$ dimensions have been proposed, none of them is known to preserve the chiral Ward identity. A given prescription then requires an additional renormalization constant or an evanescent counter-term to preserve this identity. In this letter, we use Larin’s prescription Larin (1993) and replace $/\!\!\!p_{a}\gamma_{5}$ by ${i\over 6}~{}\epsilon_{\mu\nu\sigma\lambda}p_{a}^{\mu}\gamma^{\nu}\gamma^{\sigma}\gamma^{\lambda}$. The product of two Levi-Civita tensors is computed through the determinant of Kronecker deltas defined in $D$ dimensions. The UV singularities are regulated through the renormalization of the strong coupling at the scale $\mu_{R}$. The IR singularities cancel among virtual and real emission processes, except those from either incoming or tagged final state partons that are collinear to the rest of partons. Mass factorization guarantees that the partonic cross sections in eq. (6) factorise into the spin-dependent Altarelli-Parisi (AP) kernels $\Delta\Gamma_{c\leftarrow a}$ of PDFs and $\tilde{\Gamma}_{b\leftarrow d}$ of FFs, appropriately convoluted with the finite CF (${\cal G}_{I,cd}$) at an arbitrary scale $\mu_{F}$ (suppressed here for brevity),
@@
Beyond LO, the contributions to $d\Delta\hat{\sigma}_{1,ab,L}$ can be classified into three categories: pure virtual (VV), pure real emissions (RR) and interference of real emission and virtual (RV). The VV part gets contributions from one-loop and two-loop virtual corrections to the Born process. The latter can be obtained using the quark form factor, see Lee et al. (2022). For the rest, we follow the standard Feynman diagrammatic approach. We use QGRAF Nogueira (1993) to generate Feynman diagrams and use a set of in-house routines written in FORM Kuipers et al. (2013); Ruijl et al. (2017), to convert the output of QGRAF into a suitable format to apply Feynman rules and to perform Dirac algebra, Lorentz contractions and simplifications of color factors. The computations of phase-space integrals are challenging compared to those required for inclusive cross sections because of the presence of an additional constraint $\frac{z}{z_{1}}=\frac{p_{a}\cdot p_{b}}{p_{a}\cdot q}$. The two-body phase-space over one-loop Feynman integrals that appear in RV and three-body phase space integrals in RR are simplified with reverse unitarity Anastasiou et al. (2004, 2012). This method allows us to apply loop-integration techniques, namely integration-by-parts identities (IBP) Chetyrkin and Tkachov (1981); Laporta (2000), to reduce the phase-space integrals to a smaller number of the master integrals (MIs). The constraint $\frac{z}{z_{1}}=\frac{p_{a}\cdot p_{b}}{p_{a}\cdot q}$ is introduced through the delta function $\delta\left(z^{\prime}-\frac{p_{a}\cdot p_{b}}{p_{a}\cdot q}\right)$, recall $z^{\prime}=z/z_{1}$, which is replaced by a propagator-like term $-\frac{1}{\pi}\text{Im}(1/(z^{\prime}-\frac{p_{a}\cdot p_{b}}{p_{a}\cdot q}+i\epsilon))$ with $p_{b}=p_{a}+q-k_{1}$ or $p_{b}=p_{a}+q-k_{1}-k_{2}$ for two- and three-body final states respectively. To perform the IBP reduction, we use the Mathematica package LiteRed Lee (2014).
@@
After IBP reduction, we end up with 7 MIs for RV and 20 MIs for RR sub-processes. Due to the delta function constraint, the results of the MIs depend on two scaling variables ($x^{\prime},z^{\prime}$). We have used two different approaches to compute these integrals. In the first approach, we choose a convenient Lorentz frame to parameterize the momenta so that the constraint on $z^{\prime}$ takes the simple form and three-body phase-space integrals become three-dimensional parametric integrals, see Matsuura et al. (1989); Zijlstra and van Neerven (1992); Rijken and van Neerven (1997); Ravindran et al. (2003) for more details. We encounter two angular integrals and one parametric integral. Angular integrals reduce to hypergeometric functions and the parametric integrals over these functions lead to multiple polylogarithms (MPLs) and Nielsen polylogarithms of weight up to three. In the second approach, we use the method of differential equations (DEs) Kotikov (1991); Argeri and Mastrolia (2007); Remiddi (1997); Henn (2013); Ablinger et al. (2016) to solve the integrals. We set up the system of differential equations of the MIs with respect to the variables $x^{\prime},z^{\prime}$ using LiteRed. Each set of DEs is controlled by a 20$\times$20 matrix. By an appropriate set of transformations on the set of MIs, we can express these matrices in an upper or lower-triangular form leading to the bottom-up approach of solving the DEs one by one. Alternatively, we use the elegant approach of an $\varepsilon$-factorized form Henn (2013) to reduce the DEs to canonical form with the help of the Mathematica package Libra Lee (2021). We use suitable boundary conditions to express the solution in terms of either classical polylogarithms or generalized harmonic polylogarithms (GPLs). The boundary conditions for the MIs are computed in the threshold limit from parametric integrals. We encounter four types of square-roots in the DE systems: $\big{(}\sqrt{x^{\prime}}$, $\sqrt{z^{\prime}}$, $\sqrt{(1+x^{\prime})^{2}-4x^{\prime}z^{\prime}}$, $\sqrt{(1-z^{\prime})^{2}+4x^{\prime}z^{\prime}}\big{)}$. Thanks to suitable transformations on $x^{\prime},z^{\prime}$, we can express all the polylogarithms or GPLs with simple indices, ready for numerical evaluations.
@@
×$\hskip 34.14322pt\times{}_{2}F_{1}\left(\frac{\varepsilon}{2},\frac{\varepsilon}{2};1+\frac{\varepsilon}{2};\frac{x^{\prime}(1-x^{\prime})}{(x^{\prime}-z^{\prime})(1-z^{\prime}-x^{\prime})}\right)$ (10)
@@
where, $\theta_{1a}=\theta(z^{\prime}-x^{\prime})$, $\theta_{1b}=\theta(x^{\prime}-z^{\prime})$, $\theta_{2a}=\theta(1-z^{\prime}-x^{\prime})$, $\theta_{2b}=\theta(z^{\prime}+x^{\prime}-1)$, $A$=$\Big{|}\frac{(z^{\prime}-x^{\prime})(1-z^{\prime}-x^{\prime})}{z^{\prime}(1-z^{\prime})}\Big{|}^{-\frac{\varepsilon}{2}}$, $F(\varepsilon,y)={}_{2}F_{1}(\frac{\varepsilon}{2},\frac{\varepsilon}{2};1+\frac{\varepsilon}{2};y)$ and $B=\frac{x^{\prime}(1-x^{\prime})}{(x^{\prime}-z^{\prime})(1-z^{\prime}-x^{\prime})}$. $F(\varepsilon,B\pm i\epsilon)$ can be analytically continued to the appropriate region and expanded in a power series in $\varepsilon$, see Duplancic and Nizic (2001); Gehrmann and Remiddi (2002); Gehrmann and Schürmann (2022). Finally, collinear singularities in $w=x^{\prime},z^{\prime}$ are exposed through
@@
±$\qquad\pm\Delta f_{\bar{q}}(x)D_{q^{\prime}}(z)+\Delta f_{\bar{q}}(x)D_{\bar{q}^{\prime}}(z)\,,$
@@ Acknowledgements.
Acknowledgements: We thank W. Vogelsang and S. Weinzierl for discussions. This work has been supported through a joint Indo-German research grant by the Department of Science and Technology (DST/INT/DFG/P-03/2021/dtd.12.11.21). S.M. acknowledges the ERC Advanced Grant 101095857 Conformal-EIC. The work of R.L. was supported via RSF grant No. 20-12-00205.
Note added:
While finalizing this work, ref. Bonino et al. (2024b) has appeared on polarized SIDIS at NNLO.

### v3

@@ Abstract
$\times\mathcal{G}_{1,ab}\left(\frac{x}{x_{1}},\frac{z}{z_{1}},\mu_{F}^{2},Q^{2}\right)\,$, (4)
@@
Beyond LO in perturbative QCD, we encounter both ultraviolet (UV) and infrared (IR) singularities. The latter are due to the presence of soft and collinear partons. We regulate these singularities using dimensional regularization with $D=4+\varepsilon$ space-time dimensions. The projection of spin-dependent partonic amplitudes squared $|\Delta M_{ab}|^{2}$ in eq. (6) requires Dirac matrices $\gamma_{5}$ or the Levi-Civita tensor for polarized quarks or gluons, respectively, see, e.g. Zijlstra and van Neerven (1994). Since $\gamma_{5}$ and the Levi-Civita tensor are intrinsically four-dimensional objects, their treatment in $D$ dimensions requires some prescription. Although, several schemes to define them in $D$ dimensions have been proposed, none of them is known to preserve the chiral Ward identity. A given prescription then requires an additional renormalization constant or an evanescent counter-term to preserve this identity. In this letter, we use Larin’s prescription Larin (1993) and replace $\penalty 50\!\!\!p_{a}\gamma_{5}$ by ${\frac{i}{6}}~\epsilon_{\mu\nu\sigma\lambda}p_{a}^{\mu}\gamma^{\nu}\gamma^{\sigma}\gamma^{\lambda}$. The product of two Levi-Civita tensors is computed through the determinant of Kronecker deltas defined in $D$ dimensions. The UV singularities are regulated through the renormalization of the strong coupling at the scale $\mu_{R}$. The IR singularities cancel among virtual and real emission processes, except those from either incoming or tagged final state partons that are collinear to the rest of partons. Mass factorization guarantees that the partonic cross sections in eq. (6) factorise into the spin-dependent Altarelli-Parisi (AP) kernels $\Delta\Gamma_{c\leftarrow a}$ of PDFs and $\tilde{\Gamma}_{b\leftarrow d}$ of FFs, appropriately convoluted with the finite CF (${\cal G}_{I,cd}$) at an arbitrary scale $\mu_{F}$ (suppressed here for brevity),
@@
Beyond LO, the contributions to $d\Delta\hat{\sigma}_{1,ab,L}$ can be classified into three categories: pure virtual (VV), pure real emissions (RR) and interference of real emission and virtual (RV). The VV part gets contributions from one-loop and two-loop virtual corrections to the Born process. The latter can be obtained using the quark form factor, see Lee et al. (2022). For the rest, we follow the standard Feynman diagrammatic approach. We use QGRAF Nogueira (1993) to generate Feynman diagrams and use a set of in-house routines written in FORM Kuipers et al. (2013); Ruijl et al. (2017), to convert the output of QGRAF into a suitable format to apply Feynman rules and to perform Dirac algebra, Lorentz contractions and simplifications of color factors. The computations of phase-space integrals are challenging compared to those required for inclusive cross sections because of the presence of an additional constraint $\frac{z}{z_{1}}=\frac{p_{a}\cdot p_{b}}{p_{a}\cdot q}$. The two-body phase-space over one-loop Feynman integrals that appear in RV and three-body phase space integrals in RR are simplified with reverse unitarity Anastasiou et al. (2004, 2012). This method allows us to apply loop-integration techniques, namely integration-by-parts identities (IBP) Chetyrkin and Tkachov (1981); Laporta (2000), to reduce the phase-space integrals to a smaller number of the master integrals (MIs). The constraint $\frac{z}{z_{1}}=\frac{p_{a}\cdot p_{b}}{p_{a}\cdot q}$ is introduced through the delta function $\delta\left(z^{\prime}-\frac{p_{a}\cdot p_{b}}{p_{a}\cdot q}\right)$, recall $z^{\prime}=z/z_{1}$, which is replaced by a propagator-like term $\frac{1}{\pi}\text{Im}(1/(z^{\prime}-\frac{p_{a}\cdot p_{b}}{p_{a}\cdot q}+i\epsilon))$ with $p_{b}=p_{a}+q-k_{1}$ or $p_{b}=p_{a}+q-k_{1}-k_{2}$ for two- and three-body final states respectively. To perform the IBP reduction, we use the Mathematica package LiteRed Lee (2014).
@@
After IBP reduction, we end up with 7 MIs for RV and 20 MIs for RR sub-processes. Due to the delta function constraint, the results of the MIs depend on two scaling variables ($x^{\prime},z^{\prime}$). We have used two different approaches to compute these integrals. In the first approach, we choose a convenient Lorentz frame to parameterize the momenta so that the constraint on $z^{\prime}$ takes the simple form and three-body phase-space integrals become three-dimensional parametric integrals, see Matsuura et al. (1989); Zijlstra and van Neerven (1992); Rijken and van Neerven (1997); Ravindran et al. (2003) for more details. We encounter two angular integrals and one parametric integral. Angular integrals reduce to hypergeometric functions and the parametric integrals over these functions lead to multiple polylogarithms (MPLs) and Nielsen polylogarithms of weight up to three. In the second approach, we use the method of differential equations (DEs) Kotikov (1991); Argeri and Mastrolia (2007); Remiddi (1997); Henn (2013); Ablinger et al. (2016) to solve the integrals. We set up the system of differential equations of the MIs with respect to the variables $x^{\prime},z^{\prime}$ using LiteRed. Each set of DEs is controlled by a $\times20$ matrix. By an appropriate set of transformations on the set of MIs, we can express these matrices in an upper or lower-triangular form leading to the bottom-up approach of solving the DEs one by one. Alternatively, we use the elegant approach of an $\varepsilon$-factorized form Henn (2013) to reduce the DEs to canonical form with the help of the Mathematica package Libra Lee (2021). We use suitable boundary conditions to express the solution in terms of either classical polylogarithms or generalized harmonic polylogarithms (GPLs). The boundary conditions for the MIs are computed in the threshold limit from parametric integrals. We encounter four types of square-roots in the DE systems: $\big(\sqrt{x^{\prime}}$, $\sqrt{z^{\prime}}$, $\sqrt{(1+x^{\prime})^{2}-4x^{\prime}z^{\prime}}$, $\sqrt{(1-z^{\prime})^{2}+4x^{\prime}z^{\prime}}\big)$. Thanks to suitable transformations on $x^{\prime},z^{\prime}$, we can express all the polylogarithms or GPLs with simple indices, ready for numerical evaluations.
@@
$\hskip 34.14322pt\times{}_{2}F_{1}\left(\frac{\varepsilon}{2},\frac{\varepsilon}{2};1+\frac{\varepsilon}{2};\frac{x^{\prime}(1-x^{\prime})}{(x^{\prime}-z^{\prime})(1-z^{\prime}-x^{\prime})}\right)$ (10)
@@
[…]
@@
$\qquad\pm\Delta f_{\bar{q}}(x)D_{q^{\prime}}(z)+\Delta f_{\bar{q}}(x)D_{\bar{q}^{\prime}}(z)\,$,
@@ Acknowledgements.
Acknowledgements: We thank W. Vogelsang and S. Weinzierl for discussions. We thank the authors of Bonino et al. (2024b, 2025) for correspondence regarding the finite renormalization constant $Z_{ab}$. The phenomenological results presented in the original paper are unaffected, as numerical impact of the modifications is not numerically significant. This work has been supported through a joint Indo-German research grant by the Department of Science and Technology (DST/INT/DFG/P-03/2021/dtd.12.11.21). S.M. acknowledges the ERC Advanced Grant 101095857 Conformal-EIC. The work of R.L. was supported via RSF grant No. 20-12-00205.